a large number of bullets are fired in all directions with same speed v. the maximum area on the ground on which these bullets will spread is

Grade:

4 Answers

Sunil Kumar FP
9 years ago
maximum horizotal range of aprojectile is at angle of 45 degree.
range at this angle=v^2/g=radius of the cirrcle of maximum area
therefore area=pi v^4/g^2
thanks and regards
sunil kr
askIITian faculty
naveen
6 years ago
πu square by g square is the answerWhere u is initial velocity g is acceleration due to gravityHai my name is Naveen samileti
Anumita Ghose
5 years ago
Maximum Range, R = u² sin 2 (45)/g
                            =>R = v²/g
Now, Range becomes a the radius of the circular area in which the bullets are fired.
Area = πR² = π (v²/g)² => πv⁴/g²
 
With Regards, 
Anumita Ghose
ankit singh
3 years ago
XPLANATION:
    The farthest the bullet lands will be the bullets which are launched at an angle of 45^{\circ}. And the farthest they will go will be v^{2} / g. So the radius of the circle which the bullets cover is v^{2} / g.
  Hence the area is P I \times R^{2}. So the area is\mathrm{P} I \times \mathrm{v}^{4} / \mathrm{g}^{2}.